OpenStudy (anonymous):

Solve: 10/x - 3/x-1 = 2

OpenStudy (anonymous):

start by simplifying... 10/x - 3/x = 7/x and -1 moved to the other side = 3

OpenStudy (anonymous):

so 7/x = 3 ... from there you multiply x on both sides to reach 7 = 3x... then divide the three back to the other side to get x = 7/3.

OpenStudy (anonymous):

On the paper in front of me with that questions the answer 7/3 isn't even listed but these are: a) {-6/7} b) {2,5/2} c) {12/7} d) {-2,-5/4} e) none of these

OpenStudy (anonymous):

Are you looking for the value of x?

OpenStudy (anonymous):

I have no idea what I am looking for, cause all this problem says is "SOLVE"

OpenStudy (anonymous):

I think the first guy's answer is correct. So I think you should go with "e) none of these"

OpenStudy (anonymous):

Okay. I'll go with none of these. ~ for now

OpenStudy (anonymous):

Is it - 3/x -1 or - 3 over x-1?

OpenStudy (anonymous):

The answer choices suggest a quadratic.

OpenStudy (anonymous):

b is the correct answer. http://www.wolframalpha.com/input/?i=10%2Fx+-+3%2F%28x-1%29+%3D+2